### v2

# Fast Monte Carlo Tree Diffusion: 100× Speedup via Parallel Sparse Planning

Abstract
Diffusion models have recently emerged as a powerful approach for trajectory planning. However, their inherently non-sequential nature limits their effectiveness in long-horizon reasoning tasks at test time. The recently proposed Monte Carlo Tree Diffusion (MCTD) offers a promising solution by combining diffusion with tree-based search, achieving state-of-the-art performance on complex planning problems. Despite its strengths, our analysis shows that MCTD incurs substantial computational overhead due to the sequential nature of tree search and the cost of iterative denoising. To address this, we propose Fast-MCTD, a more efficient variant that preserves the strengths of MCTD while significantly improving its speed and scalability. Fast-MCTD integrates two techniques: Parallel MCTD, which enables parallel rollouts via delayed tree updates and redundancy-aware selection; and Sparse MCTD, which reduces rollout length through trajectory coarsening. Experiments show that Fast-MCTD achieves up to 100× speedup over standard MCTD while maintaining or improving planning performance. Remarkably, it even outperforms Diffuser in inference speed on some tasks, despite Diffuser requiring no search and yielding weaker solutions. These results position Fast-MCTD as a practical and scalable solution for diffusion-based inference-time reasoning.
1 Introduction
Diffusion models have recently emerged as a powerful paradigm for trajectory planning, leveraging offline datasets to generate complex, high-quality trajectories through iterative denoising [12, 2, 34, 1]. Unlike autoregressive planners that generate trajectories sequentially via forward dynamics [10, 11, 3], diffusion-based approaches like Diffuser [12] generate trajectories holistically, mitigating issues such as long-term dependencies and cumulative error.
Despite these advantages, diffusion planners often struggle with complex reasoning at test time, especially in long-horizon tasks. They may produce plausible but suboptimal trajectories that fail to accomplish complex goals. To address this, inference-time scaling has emerged as a promising approach to enhance reasoning by adding computational procedures—such as verification and refinement—during inference [34, 33, 20, 19, 25, 15].
However, designing a diffusion-based method that supports inference-time scaling remains challenging. This is primarily due to a fundamental mismatch: while reasoning often requires sequential or causal processing, diffusion models are inherently non-sequential and generate trajectories in a non-causal manner by design. As a result, simple strategies such as increasing the denoising depth or employing best-of-$N$ sampling have been shown to offer only limited improvements [31].
To address this challenge, Yoon et al. [31] proposed Monte Carlo Tree Diffusion (MCTD), which integrates diffusion-based planning with sequential search through tree-based reasoning, akin to Monte Carlo Tree Search (MCTS) [7]. MCTD adopts Diffusion Forcing [2] as the backbone and reinterprets its block-wise denoising as a causal tree rollout, introducing sequential structure while preserving the global generative strengths of diffusion models. By structuring trajectory generation as a tree search, MCTD enables systematic exploration and exploitation, helping the model escape local optima and discover higher-quality trajectories.
While MCTD has demonstrated impressive performance in complex, long-horizon tasks that conventional methods fail to solve, further analysis reveals that it suffers from significant computational inefficiencies due to two key bottlenecks: the sequential nature of MCTS, which updates tree statistics after each iteration, and the iterative denoising process inherent to diffusion models. Unfortunately, this inefficiency is most pronounced in the very long-horizon settings where MCTD’s planning capabilities are most beneficial. Thus, improving the efficiency of MCTD is the most critical challenge for establishing it as a broadly practical solution for diffusion-based inference-time scaling.
In this paper, we propose Fast Monte Carlo Tree Diffusion (Fast-MCTD), a framework that significantly reduces the computational overhead of tree search and iterative denoising while preserving the strong planning capabilities of MCTD. Fast-MCTD integrates two key optimization techniques: Parallel MCTD (P-MCTD) and Sparse MCTD (S-MCTD). P-MCTD accelerates the tree search process by enabling parallel rollouts, deferring tree updates until multiple searches are completed, introducing redundancy-aware selection, and parallelizing both expansion and simulation steps. S-MCTD further improves efficiency by planning over coarsened trajectories, using diffusion models trained on these compressed representations. This not only reduces the cost of iterative denoising but also lowers overall search complexity by effectively shortening the planning horizon.
Experimental results show that Fast-MCTD achieves substantial speedups—up to 100× on some tasks—compared to standard MCTD, while maintaining comparable or superior planning performance. Remarkably, Fast-MCTD also outperforms Diffuser in inference speed on some tasks, despite Diffuser’s lack of search and its substantially inferior performance.
The main contributions of this paper are twofold. First, we introduce Fast-MCTD, a framework that improves the efficiency of MCTD through parallelization and rollout sparsification. Second, we empirically show that Fast-MCTD achieves substantial speedups—up to two orders of magnitude—while maintaining strong planning performance, demonstrating its practical effectiveness in challenging long-horizon tasks.
2 Preliminaries
2.1 Diffusion models for planning
Diffuser [12] formulates planning as a generative denoising process over full trajectories, defined as
$\mathbf{x}=\begin{bmatrix}s_{0}&s_{1}&\cdots&s_{T}\\ a_{0}&a_{1}&\cdots&a_{T}\end{bmatrix},$ (1)
where $T$ is the trajectory length and $(s_{t},a_{t})$ denotes the state and action at time $t$, respectively. During inference, trajectories are iteratively denoised from noisy samples, effectively reversing a forward diffusion process. Since the learned denoising model alone does not inherently optimize for returns or task objectives, Diffuser incorporates a guidance function $\mathcal{J}_{\phi}(\mathbf{x})$ inspired by classifier-guided diffusion [8]. Specifically, the guidance function biases trajectory generation towards high-return outcomes by modifying the sampling distribution as:
$\tilde{p}_{\theta}(\mathbf{x})\propto p_{\theta}(\mathbf{x})\exp\left(\mathcal{J}_{\phi}(\mathbf{x})\right),$ (2)
thus explicitly guiding the denoising process toward optimal trajectories at test time.
Diffusion Forcing [2] extends the Diffuser framework by introducing token-level denoising control within trajectories. Specifically, Diffusion Forcing tokenizes the trajectory $\mathbf{x}$, enabling different tokens to be denoised at distinct noise levels. This selective, partial denoising allows the model to generate only the tokens exhibiting higher uncertainty, such as future plan tokens.
2.2 MCTD: Monte Carlo Tree Diffusion
MCTD unifies tree search and diffusion-based planning by integrating three key concepts.
(1) Denoising as tree rollout. Unlike traditional MCTS, which expands trees over individual states, MCTD partitions a trajectory into temporally extended subplans ${\bf x}=({\bf x}_{1},\dots,{\bf x}_{S})$, each treated as a high-level node. It applies a semi-autoregressive denoising schedule using Diffusion Forcing, where earlier subplans are denoised faster, conditioning later ones on previous outputs, approximating:
where $g_{s}$ is a guidance level (weight) for the classifier-guided sampling [8] (Equation 2). This preserves the global coherence of diffusion models while enabling intermediate rollouts akin to MCTS, substantially reducing tree depth.
(2) Guidance levels as meta-actions. To address exploration-exploitation trade-offs in large or continuous action spaces, MCTD introduces meta-actions via guidance schedules. A guidance schedule ${\bf g}=(g_{1},\dots,g_{S})$ assigns discrete control modes—e.g., guide or no_guide—to each subplan, modulating whether it is sampled from an unconditional prior or a guided distribution (Equation 2). This allows selective guidance within a single diffusion process, enabling flexible trade-off control.
(3) Jumpy Denoising as fast simulation. For efficient simulation, MCTD uses jumpy denoising (e.g., Denoising Diffusion Implicit Models (DDIM) [27]) to skip denoising steps and rapidly complete remaining trajectory segments. Given subplans ${\bf x}_{1:s}$, the rest $\tilde{\bf x}_{s+1:S}$ is coarsely denoised as: $\tilde{\bf x}_{s+1:S}\sim p({\bf x}_{s+1:S}|{\bf x}_{1:s},{\bf g})$, yielding a full trajectory $\tilde{\bf x}$ for evaluation.
The four steps of an MCTD round. These components are instantiated through the four canonical steps of MCTS adapted to operate over subplans within a diffusion-based planning framework. See [31] for a detailed pseudocode of the MCTD method.
  1. 1.
    Selection. MCTD traverses the tree from the root to a leaf using a selection strategy like Upper Confidence Boundary of Tree (UCT) [14]. Each node corresponds to a temporally extended subplan, not a single state, reducing tree depth and improving abstraction. The guidance schedule ${\bf g}$ is updated during traversal to balance exploration (no_guide) and exploitation (guide).
  2. 2.
    Expansion. Upon reaching a leaf, a new subplan ${\bf x}_{s}$ is sampled using the diffusion forcing, conditioned on the chosen meta-action $g_{s}$. Sampling may follow an exploratory prior $p({\bf x}_{s}|{\bf x}_{1:s-1})$ or a goal-directed distribution $p({\bf x}_{s}|{\bf x}_{1:s-1},g_{s})$.
  3. 3.
    Simulation. Fast jumpy denoising (e.g., via DDIM) completes the remaining trajectory $\tilde{\bf x}$ after expansion. Although approximate, this is computationally efficient and sufficient for plan evaluation using a reward function $r(\tilde{\bf x})$.
  4. 4.
    Backpropagation. The obtained reward is backpropagated through the tree to update value estimates and refine guidance strategies for future planning rounds. This enables adaptive, reward-informed control of exploration and exploitation.
3 The Planning Horizon Dilemma
Despite its strong performance on long-horizon and complex planning tasks, MCTD suffers from a fundamental computational inefficiency. This is because, it requires repeated partial denoising operations to generate feasible subplans. Worse yet, as the planning horizon increases, the search space grows exponentially, resulting in significantly higher computational cost.
Ironically, the key strength of diffusion-based planning—its capacity for long-horizon reasoning—becomes a principal bottleneck when integrated with the inherently sequential nature of Monte Carlo Tree Search. In effect, the very scenarios where MCTD’s planning capabilities are most advantageous are also those where it incurs the highest computational cost. We refer to this tension between the global trajectory reasoning of Diffuser and the step-by-step search process of MCTS as the Planning Horizon Dilemma.
More formally, consider a diffusion forcing schedule that generates a sequence of subplans. Let $N_{\text{child}}$ denote the branching factor at each tree node, corresponding to the cardinality of the meta-action space. Let $C_{\text{sub}}$ be the computational cost of performing partial denoising for a single subplan ${\bf x}_{s}$. Lastly, let $\bar{s}\leq S$ be the number of subplans required to reach the goal. Then, the total complexity of identifying a successful trajectory using MCTD is given by:
${C_{\text{MCTD}}={\cal O}(N_{\text{child}}^{\bar{s}}\cdot C_{\text{sub}})\,.}$ (4)
Although MCTD’s exploration-exploitation strategy reduces average-case complexity, its computational cost still grows exponentially with the expected subplan length $\bar{s}$.
This inefficiency of MCTD is clearly demonstrated in Figure 1, which compares planning time and success rates across two maze environments (point mass and ant robot) and two map sizes (medium and giant). As maze size increases from medium to giant, other diffusion-based planners such as Diffuser [12] and Diffusion Forcing [2] exhibit substantial drops in performance. In contrast, MCTD maintains a significant performance advantage due to its strong search capability. However, this advantage comes at a steep computational cost: in medium-sized mazes, MCTD requires roughly 8–10× more planning time than Diffuser, which further increases to 30–40× in giant-sized environments.
Figure 1: Planning time vs. success rate. As maze size increases, most diffusion-based planners degrade in performance. MCTD maintains high success rates but with long planning times, reflecting the Planning Horizon Dilemma—better performance requires more computation. Fast-MCTD breaks this trade-off, achieving strong performance with much faster planning.
4 Fast Monte Carlo Tree Diffusion
The Planning Horizon Dilemma reveals two primary sources of inefficiency in MCTD: (1) Between-rollout inefficiency — partial denoising operations across different rollouts are executed serially, limiting parallelism; and (2) Within-rollout inefficiency — each rollout requires the Diffusion to process long trajectories, leading to substantial computational overhead. To address these challenges, we propose Fast-MCTD, which incorporates two key improvements: Parallel Planning via P-MCTD, to improve concurrency, and Sparse Planning via S-MCTD, to reduce the effective planning horizon. The overall processes are illustrated in Figure 2.
4.1 Parallel planning
Independently Parallel MCTD via Delayed Tree Update. We begin by parallelizing all four stages of MCTD through $K$ concurrent rollouts. In this setting, each iteration generates a batch of tree rollouts, each proceeding independently within the shared search tree. To avoid synchronization overhead, we adopt a delayed tree update strategy: all rollouts operate on a shared, fixed snapshot of the tree, and updates to the tree (e.g., value estimates and visitation counts) are applied only after all rollouts in the batch are completed.
While this design allows for efficient parallel execution, it introduces a trade-off. As the batch size increases, the tree statistics used during search become increasingly stale, which can degrade planning performance by reducing the accuracy of selection and value propagation. To address this inefficiency, we propose a mitigation strategy in the following section. We further analyze this trade-off empirically in Section 6.4, identifying the optimal degree of parallelism that balances computational efficiency and search quality.
Redundancy-Aware Selection. While the delayed tree update enables parallel rollouts, it introduces a critical inefficiency: redundant node selection. That is, multiple rollouts within a batch independently select and thus may expand the same node, resulting in duplicated computation and limited exploration diversity. More specifically, this issue arises because the standard node selection policy in MCTD—Upper Confidence Bound for Trees (UCT) [14]—is designed for sequential search and does not account for concurrent expansions.
Table 1: Each step time of MCTD on PointMaze-Giant
MCTD Step Time (sec.)
Selection 3e-4 (0.05%)
Expansion 0.393 (70.9%)
Simulation 0.161 (29.0%)
Backpropagation 1e-4 (0.02%)
To address this issue, we introduce the Redundancy-Aware Selection (RAS) mechanism. Our key observation motivating this design is that the time spent on the Selection step, among the four stages of MCTD, is negligibly small: as shown in Table 1, it only takes 0.05% time of the MCTD steps. This allows us to remove the Selection (and the Backpropagation) step from the parallelization pipeline and instead implement it as a lightweight serial operation.
Specifically, inspired by prior work on parallelized MCTS [24, 16, 18], we implement Redundancy-Aware Selection (RAS) by modifying the standard UCT-based selection criterion. During each parallel search phase, we temporarily incorporate an auxiliary visitation count variable $\smash{{\hat{N}_{i}}}$, which sequentially tracks selections made within the current batch. This variable is reset to zero after the delayed tree update is applied. The resulting selection policy is defined as follows:
${\pi(i)=\operatorname*{arg\,max}_{j\in\mathcal{C}(i)}\left(V_{j}+\beta\sqrt{\frac{\log(N_{i}+\hat{N}_{i}\cdot w)}{N_{j}+\hat{N}_{j}\cdot w}}\right),}$ (5)
where $\pi(i)$ denotes the node selection policy from node $i$, $\mathcal{C}(i)$ is the set of child nodes of node $i$, and $V_{i}$, $N_{i}$ represent the estimated value and visitation count of node $i$, respectively. The hyperparameter $w$ adjusts the exploration-exploitation balance for parallel search: when $w=0$, the policy reduces to standard MCTD selection, and higher values of $w$ penalize nodes that have already been selected during the current batch, encouraging other rollouts to explore different parts of the tree.
Parallel Denoising on Expansion and Simulation. Unlike conventional MCTS, where computational costs primarily stem from simulations, MCTD incurs substantial overhead in both expansion and simulation due to the expensive denoising operations. To enhance computational efficiency, we propose a unified batching strategy, termed parallel denoising, that simultaneously processes multiple subplans chosen by RAS during the expansion and simulation phases. Specifically, we implement a common parallel_subplan interface that batches denoising steps by scheduling noise levels and synchronizing DDIM [27] updates. To handle variable-length subplans and different guidance levels, subplans are zero-padded and packed into uniformly shaped tensors, enabling high-throughput parallel execution on GPUs.
(a) Parallel MCTD
(b) Sparse MCTD
Figure 2: Two key components of Fast-MCTD. (a) Parallel MCTD accelerates planning by performing batched expansion and simulation on a partial denoising tree, followed by delayed tree updates. (b) Sparse MCTD reduces denoising overhead by planning over abstract sub-trajectories, significantly decreasing the number of subplans.
4.2 Sparse planning
Another key contributor to the Planning Horizon Dilemma is the within-rollout inefficiency, wherein the cost of denoising increases with length of the rollout trajectories. This inefficiency persists even with the parallelization improvements introduced by P-MCTD. To address this issue, we propose Sparse Monte Carlo Tree Diffusion (S-MCTD). The key idea is to incorporate trajectory coarsening—operating rollouts at a higher level of abstraction, as suggested in prior work [5, 4]—into the MCTD framework, thereby reducing the effective rollout length and overall computational cost.
Algorithm 1 Fast-MCTD
1:Initialize tree $\mathcal{T}$, selected node set ${\cal S}$
2:while not goal reached do
3:    Reset $\hat{N}_{i}$ for all node $i$ in ${\cal T}$
4:    for $i=1$ to $K$ do
5:        Select node $v_{i}$ from $\mathcal{T}$ Equation (5)
6:        Update ${\cal S}$: ${\cal S}={\cal S}\cup v_{i}$
7:    end for
8:    for $i=1\in{\cal S}$ in parallel do
9:        Expand node $v_{i}$ via SparseExpand
10:        Simulate with SparseJumpySimulate
11:    end for
12:    for $i=1$ to $K$ do
13:        Backpropagate rewards through $\mathcal{T}$
14:    end for
15:end while
16:return best node from $\mathcal{T}$
Specifically, prior to training the diffusion model, we construct a dataset of coarse trajectories by subsampling every $H$ steps from the original trajectories, yielding $\mathbf{x\textquoteright}=[x_{1},x_{H+1},x_{2H+1},\dots]$. These coarse trajectories are then modeled using a specialized sparse diffusion planner. Owing to the shorter trajectory lengths, the computational cost of denoising each coarse subplan—denoted as $C_{\text{coarse}}$—is significantly lower than that of the original fine-grained subplans, $C_{\text{sub}}$. Furthermore, the overall search complexity is reduced, as planning now involves approximately $H$ times fewer subplans, i.e., $\mathbf{x\textquoteright}=[\mathbf{x\textquoteright}_{1},\mathbf{x\textquoteright}_{2},\dots,\mathbf{x\textquoteright}_{S\textquoteright}]$, where $S\textquoteright\approx S/H$ given the original trajectory length $S$, as illustrated in Figure 2(b). In terms of computational complexity, S-MCTD offers a substantial reduction compared to MCTD (Equation (4)):
${C_{\text{S-MCTD}}=\mathcal{O}\left(N_{\text{child}}^{S/H}\cdot C_{\text{coarse}}\right)\,.}$ (6)
We observe that computational efficiency improves exponentially with the interval size $H$. However, excessively large intervals may overly abstract the planning task, potentially leading to degraded performance. We empirically investigate this trade-off in Section 6.4.
Following prior work [12, 4, 31], we implement low-level control for executing each coarse subplan ${\bf x}^{\prime}_{s}$ using heuristic controllers [12], value-based policies [4, 31], and inverse dynamics models [31].
4.3 Integrating Parallel and Sparse MCTD
Finally, we integrate the aforementioned Parallel MCTD (P-MCTD; Section 4.1) and Sparse MCTD (S-MCTD; Section 4.2) approaches into our final proposed method, termed Fast Monte Carlo Tree Diffusion (Fast-MCTD). We illustrate the details of Fast-MCTD in Algorithm 1.
5 Related works
Diffusion models [26] have demonstrated significant success in long-horizon trajectory planning, particularly in sparse-reward settings, by learning to generate plan holistically rather than forward modeling [12, 2, 1, 34, 17]. Since diffusion denoising is computationally intensive, many works address this limitation through hierarchical planning [5, 9, 6]. Chen et al. [4] tackles the limitation of diffusion planners in generating detailed short-term actions by incorporating value learning policies [28] for low-level control. Chen et al. [2] introduces causal noise scheduling with semi-autoregressive planning to improve performance in causality-sensitive tasks. Recent work has explored leveraging increased computational budgets to improve plan quality [31, 33]. Both Yoon et al. [31] and Zhang et al. [33] propose tree search methods for the denoising process. Yoon et al. [31] combines auto-regressive denoising with guidance scale-based action control for complex long-horizon planning, while Zhang et al. [33] develops a tree search method with learnable energy functions for value estimation, applying it to reasoning tasks such as Sudoku. However, as inference-time scaling in diffusion denoising has only recently been explored, optimizing efficiency in this context remains relatively understudied.
MCTS [7] has achieved notable success across various decision-making domains, particularly when integrated with learned policies or value networks [24, 23]. It has also been applied to Large Language Models (LLMs) to enhance reasoning required task performances [29, 30, 32]. To address MCTS’s inherent inefficiency in sequential search, various enhancements have been proposed, including parallel search methods [18], novel node selection policies [16], and visited-node penalization in parallel searches [24].
6 Experiments
We evaluate Fast-MCTD using tasks from the Offline Goal-conditioned RL benchmark (OGBench) [21], aligning with the setup in MCTD [31]. They includes tasks such as point and ant maze navigation with long horizons, robot arm multi-cube manipulation, and visually and partially observable mazes. Metrics reported are mean success rate (%) and planning time (sec), averaged over 50 runs (5 tasks $\times$ 10 seeds), including mean ± standard deviation. Detailed configurations and control settings are available in Appendix A.
Baselines. Besides MCTD, we include Diffuser [12], its replanning variant Diffuser-Replan, and Diffusion Forcing [2], which uses the causal noise schedule like MCTD but without explicit search.
Table 2: Maze results. Success rates and planning times ($\pm$ std) across PointMaze and AntMaze environments on medium, large giant sized map for navigate datasets. Best results are shown in strong positive colors, comparable ones in mild positives. For planning time, slowest results use strong negatives, and marginally slower ones use mild negatives.
Success Rate $\uparrow$ (%) Planning Time $\downarrow$ (sec.)
Env. Method medium large giant medium large giant
PointMaze Diffuser 58 $\pm$ 6 44 $\pm$ 8 0 $\pm$ 0 6.6 $\pm$ 0.1 6.5 $\pm$ 0.1 6.4 $\pm$ 0.1
Diffuser-Replan 60 $\pm$ 0 40 $\pm$ 0 0 $\pm$ 0 40.6 $\pm$ 2.7 55.1 $\pm$ 1.8 137.8 $\pm$ 0.4
Diffusion Forcing 65 $\pm$ 16 74 $\pm$ 9 50 $\pm$ 10 9.8 $\pm$ 1.0 12.5 $\pm$ 0.5 45.6 $\pm$ 3.2
MCTD 100 $\pm$ 0 98 $\pm$ 6 100 $\pm$ 0 59.2 $\pm$ 27.1 174.6 $\pm$ 27.2 264.2 $\pm$ 33.8
P-MCTD (Ours) 96 $\pm$ 8 100 $\pm$ 0 100 $\pm$ 0 7.4 $\pm$ 1.1 7.0 $\pm$ 0.6 9.8 $\pm$ 1.3
S-MCTD (Ours) 100 $\pm$ 0 84 $\pm$ 8 100 $\pm$ 0 9.7 $\pm$ 1.2 226.1 $\pm$ 114.4 18.9 $\pm$ 2.8
Fast-MCTD (Ours) 100 $\pm$ 0 80 $\pm$ 0 100 $\pm$ 0 3.6 $\pm$ 0.6 4.1 $\pm$ 0.4 2.4 $\pm$ 0.1
AntMaze Diffuser 36 $\pm$ 15 14 $\pm$ 16 0 $\pm$ 0 6.2 $\pm$ 0.1 6.5 $\pm$ 0.1 6.5 $\pm$ 0.1
Diffuser-Replan 40 $\pm$ 18 26 $\pm$ 13 0 $\pm$ 0 60.3 $\pm$ 4.2 73.7 $\pm$ 1.9 193.0 $\pm$ 0.4
Diffusion Forcing 90 $\pm$ 10 57 $\pm$ 6 24 $\pm$ 12 14.0 $\pm$ 3.8 24.1 $\pm$ 1.8 67.0 $\pm$ 3.4
MCTD 100 $\pm$ 0 98 $\pm$ 6 94 $\pm$ 9 68.0 $\pm$ 3.8 132.1 $\pm$ 12.7 214.5 $\pm$ 11.0
P-MCTD (Ours) 96 $\pm$ 8 90 $\pm$ 10 89 $\pm$ 16 2.8 $\pm$ 0.2 3.1 $\pm$ 0.1 4.2 $\pm$ 0.1
S-MCTD (Ours) 94 $\pm$ 1 68 $\pm$ 13 82 $\pm$ 14 12.5 $\pm$ 0.7 40.1 $\pm$ 2.6 40.3 $\pm$ 2.9
Fast-MCTD (Ours) 98 $\pm$ 6 77 $\pm$ 14 90 $\pm$ 16 2.2 $\pm$ 0.1 2.2 $\pm$ 0.1 2.5 $\pm$ 0.1
6.1 Long-horizon planning in maze
We evaluate our methods’ efficiency in extensive horizon scenarios using PointMaze and AntMaze tasks from OGBench [21], requiring up to 1000-step trajectories in medium, large, and giant mazes. Following  [31], PointMaze employs a heuristic controller [12], and AntMaze uses a learned value-based policy [28]. Results in Table 2 highlight that Diffuser and Diffusion Forcing exhibit performance degradation as maze sizes increase, while MCTD maintains effectiveness but at significantly increased computational costs, illustrating the planning horizon dilemma clearly. For instance, MCTD exceeds 4 minutes to solve PointMaze-Giant.
In contrast, Fast-MCTD demonstrates approximately 80-110× speedups relative to MCTD on giant maps with minimal performance loss. Surprisingly, Fast-MCTD is even faster than Diffuser, which employs straightforward denoising without search. Despite generally superior performance, Fast-MCTD encounters performance degradation on large maps due to sparse diffusion model limitations arising from insufficient start-position training data when subsampled every $H$ steps. Nevertheless, Sparse MCTD (S-MCTD) and Parallel MCTD (P-MCTD) achieve notable efficiency gains, in particular, P-MCTD reaches a 50× speedup on AntMaze-Giant.
Table 3: Robot task results. Success rates and planning times ($\pm$ std) across increasing task difficulty.
Success Rate $\uparrow$ (%) Planning Time $\downarrow$ (sec.)
Method Single Double Triple Single Double Triple
Diffuser 78 $\pm$ 23 12 $\pm$ 10 8 $\pm$ 10 6.3 $\pm$ 0.1 6.4 $\pm$ 0.1 6.5 $\pm$ 0.1
Diffuser-Replan 92 $\pm$ 13 12 $\pm$ 13 4 $\pm$ 8 21.4 $\pm$ 2.4 71.2 $\pm$ 5.1 130.9 $\pm$ 11.6
Diffusion Forcing 100 $\pm$ 0 18 $\pm$ 11 16 $\pm$ 8 2.9 $\pm$ 0.2 15.2 $\pm$ 2.2 15.9 $\pm$ 1.2
MCTD-Replan 100 $\pm$ 0 78 $\pm$ 11 40 $\pm$ 21 9.2 $\pm$ 0.9 38.8 $\pm$ 4.7 102.0 $\pm$ 6.9
P-MCTD-Replan (Ours) 100 $\pm$ 0 80 $\pm$ 9 50 $\pm$ 21 2.9 $\pm$ 0.2 6.5 $\pm$ 2.3 11.4 $\pm$ 2.8
S-MCTD-Replan (Ours) 100 $\pm$ 0 75 $\pm$ 3 42 $\pm$ 11 8.5 $\pm$ 0.5 35.7 $\pm$ 18.2 58.7 $\pm$ 33.0
Fast-MCTD-Replan (Ours) 100 $\pm$ 0 77 $\pm$ 11 50 $\pm$ 16 3.0 $\pm$ 0.2 5.9 $\pm$ 1.7 9.1 $\pm$ 1.8
6.2 Robot arm manipulation
We evaluate efficiency of our methods in compositional planning tasks involving multiple cube manipulations from OGBench [21]. Tasks require strategic manipulation sequences, such as stacking cubes in a specific order. High-level planning utilizes diffusion planners, and local actions are executed via value-learning policies [28], enhanced by object-wise guidance and replanning methods following [31].
Fast-MCTD maintains or exceeds MCTD performance with significantly improved efficiency. Specifically, it shows better efficiency than Diffuser on single and double cube tasks. However, in triple cube tasks, Fast-MCTD is slower than Diffuser due to difficulties in identifying viable plans within the limited computational budget, yet it remains approximately 10× faster than MCTD.
Table 4: Visual maze results. Mean success rate (%) and planning time (sec) on medium and large mazes.
Success Rate $\uparrow$ (%) Planning Time $\downarrow$ (sec.)
Method Medium Large Medium Large
Diffuser 8 $\pm$ 13 0 $\pm$ 0 7.0 $\pm$ 0.3 6.7 $\pm$ 0.3
Diffuser-Replan 8 $\pm$ 10 0 $\pm$ 0 26.4 $\pm$ 1.0 26.0 $\pm$ 0.4
Diffusion Forcing 66 $\pm$ 32 8 $\pm$ 12 14.5 $\pm$ 1.7 17.1 $\pm$ 0.7
MCTD 76 $\pm$ 20 2 $\pm$ 6 98.8 $\pm$ 36.2 320.1 $\pm$ 7.6
P-MCTD (Ours) 86 $\pm$ 16 0 $\pm$ 0 6.8 $\pm$ 0.4 7.6 $\pm$ 0.1
S-MCTD (Ours) 82 $\pm$ 14 31 $\pm$ 16 40.1 $\pm$ 32.0 168.7 $\pm$ 46.2
Fast-MCTD (Ours) 80 $\pm$ 18 32 $\pm$ 19 4.0 $\pm$ 0.2 5.1 $\pm$ 0.3
MCTD-Replan 86 $\pm$ 13 31 $\pm$ 10 129.8 $\pm$ 53.1 419.7 $\pm$ 49.1
Fast-MCTD-Replan (Ours) 84 $\pm$ 28 38 $\pm$ 30 14.3 $\pm$ 2.3 26.7 $\pm$ 3.7
6.3 Visual planning
To assess efficiency in high-dimensional, partially observable contexts, we evaluate Fast-MCTD using visual maze environments where agents observe RGB image pairs depicting start and goal states [31]. Observations are encoded with a pretrained VAE [13], actions derived via inverse dynamics models in latent space, and approximate positional hints are provided by pretrained estimators.
Table 4 reveals that Fast-MCTD offers a substantial efficiency improvement (25-60× faster than MCTD), even outperforming it on larger mazes. The abstract trajectory planning through Sparse MCTD (S-MCTD) reduces rollout horizon complexity and improves credit assignment in latent transitions [17] as shown in Figure 3. Consequently, S-MCTD surpasses both MCTD and P-MCTD performance on larger maps, and Fast-MCTD’s integration of both sparse and parallel strategies maximizes efficiency. Replanning further enhances MCTD but at significant computational cost, whereas Fast-MCTD achieves superior performance with relatively minimal overhead, comparable to Diffuser-Replan.
Figure 3: Planning visualization. Sparse planning (top row) allows for more effective long-horizon trajectories than the dense (bottom row).
6.4 Ablation studies
Redundancy-Unaware Selection. To assess the impact of Redundancy-Aware Selection (RAS) on planning efficiency, we ablate the constraint that prevents multiple rollouts from expanding the same node simultaneously. In this redundancy-unaware selection variant, rollouts still share visitation counts but can repeatedly expand the same node. Table 6 presents planning times of P-MCTD across different PointMaze sizes, measured under a generous computational budget with 100% success rates. By avoiding repeated expansion of the same node, RAS prevents excessive exploitation, which can dominate exploration when paths are harder to discover, especially in larger environments. In smaller environments, simple exploitation can still be effective, making the benefit of RAS less pronounced.
Visitation count weight. We further investigate the impact of RAS during the selection phase by ablating the visitation count weight w from Equation (5), which balances exploration and exploitation. Table 6 presents the planning times in the Giant PointMaze for different w values. In the standard UCT setting (w = 0), which lacks visitation count sharing, planning times exceed 220 seconds with high variance. In contrast, with RAS, planning times remain consistently low (12-14 seconds) across a wide range of $w$ values, suggesting that redundancy awareness is more critical than fine-tuning this hyperparameter.
Table 5: Planning times for Redundancy-Aware vs. Unaware Selection in PointMaze.
Planning Time $\downarrow$ (sec.)
Redundancy Medium Large Giant
Aware 8.5 $\pm$ 1.6 7.4 $\pm$ 1.3 12.8 $\pm$ 2.5
Unaware 6.6 $\pm$ 1.0 7.8 $\pm$ 0.6 18.4 $\pm$ 0.9
Table 6: Planning time with different weight $w$ in PointMaze-Giant.
Weight ($w$) Planning Time $\downarrow$ (sec.)
0.0 221.3 $\pm$ 109.6
0.1 14.0 $\pm$ 2.5
1.0 12.8 $\pm$ 2.3
5.0 12.5 $\pm$ 2.1
Parallelism degree and interval size. Here, we study two key hyperparameters: the parallelism degree (number of concurrent rollouts) and the interval size (steps skipped between generated states). Figure 4 shows the performance of Fast-MCTD in different configurations. Increasing the parallelism degree initially reduces planning time through efficient parallelization, but high degrees (above 200) degrade success rates due to delayed tree updates and selection redundancy. For interval size, larger interval sizes significantly boost efficiency but risk performance loss if the abstractions become too coarse for accurate low-level control [31].
Figure 4: Ablation studies for parallelism degree and interval size. Success rates (%) and planning time (sec.) as the parallelism degree and interval size increase for maze giant tasks.
7 Discussion
While our Fast Monte Carlo Tree Diffusion framework significantly improves performance, certain trade-offs exist. P-MCTD requires high-performance computing resources for optimal parallelization, but even with modest parallel capabilities, substantial efficiency gains are achievable on standard hardware. Our approach introduces additional hyperparameters (parallelism degree and interval size), but our ablation studies demonstrate these parameters are robust across a wide range of values with clear guidelines for effective settings. The computational benefits of Fast-MCTD substantially outweigh the minor overhead of parameter tuning, enabling practical deployment of diffusion-based planning in time-sensitive applications that were previously infeasible.
8 Conclusion
We introduced Fast Monte Carlo Tree Diffusion (Fast-MCTD), a scalable diffusion-based planning algorithm that resolves the planning horizon dilemma inherent to MCTD. By combining parallelized tree search (P-MCTD) with sparse sub-trajectory abstraction (S-MCTD), Fast-MCTD achieves significant inference-time speedups—up to 100× faster than MCTD in some tasks, while maintaining or exceeding performance across challenging long-horizon tasks. We demonstrated its effectiveness in maze navigation, robotic manipulation, and visual planning benchmarks. These gains are enabled by algorithmic innovations such as search-aware parallel rollouts and coarse-to-fine diffusion over abstract plans. While Fast-MCTD alleviates core efficiency bottlenecks, further improvements may be possible by integrating adaptive sparsity or learning-based guidance selection. Our findings highlight that structured reasoning from test-time scalability and efficient inference process are not mutually exclusive, opening new avenues for fast, deliberative decision-making in high-dimensional domains.
Acknowledgments and Disclosure of Funding
This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT. YB acknowledges the support from CIFAR, NSERC, the Future of Life Institute.
Appendix A Experiment details
A.1 Computation resources
The experiments were conducted on a computing server equipped with 8 NVIDIA RTX 4090 GPUs, 512GB system memory, and a 96-thread CPU. Model training required approximately 3-6 hours per model, while inference for each experimental evaluation took approximately up to 5 minutes.
(a) Maze Environment
(b) Robot Cube
(c) Visual Maze
Figure 5: The task illustrations: (a) long-horizon maze navigation, (b) multi-object robotic manipulation requiring compositional planning, and (c) visual maze planning from high-dimensional raw RGB observations.
A.2 Environment details
Following prior work [31], we evaluated our methods (P-MCTD, S-MCTD and Fast-MCTD) on challenging scenarios from the goal-conditioned task benchmark, OGBench [21], including long-horizon maze navigation, robot arm manipulation, and visually observable maze environments as illustrated in Figure 4. All evaluation tasks were unseen during training, requiring the models to generalize their learned planning capabilities to novel scenarios at inference time.
A.2.1 Long-horizon maze
To assess performance on extensive planning horizons, we tested our methods on point-mass and ant robot navigation tasks across three increasingly complex environments: medium, large, and giant mazes. Models were trained on the navigate dataset containing long trajectories but without specific task knowledge. At inference time, models must leverage their learned capabilities to solve previously unseen maze configurations, testing their generalization to complex long-horizon planning problems.
A.2.2 Robot arm manipulation
The robot manipulation tasks require compositional planning to move multiple objects in specific sequences. For example, when tasked with stacking objects in a predetermined order, the model must generate a coherent plan that respects these sequential constraints. As demonstrated in prior work [31], holistic diffusion-based planners often struggle with such compositional tasks, particularly without explicit search or reward guidance. We evaluated Fast-MCTD in this challenging domain, focusing on efficiency improvements within a replanning framework while maintaining MCTD’s strong performance on compositional tasks.
A.2.3 Visual maze
To evaluate MCTD in high-dimensional, partially observable settings, previous work [31] utilized the Visual Antmaze task from OGBench [21]. In this task, the agent must navigate from an initial observation to a target, both represented as 64×64 RGB images with 3 color channels. The images are encoded into an 8-dimensional latent space using a Variational Autoencoder (VAE) [13], providing a compact representation for efficient planning.
A.3 Baselines
We compared Fast-MCTD against several diffusion-based planning approaches:
  • Diffuser [12]: We implemented the standard Diffuser with classifier-guided generation [8] as our foundational baseline [31], providing comparable guidance mechanisms across all tested methods.
  • Diffuser-Replan: To strengthen this baseline, we extended Diffuser with periodic replanning capabilities. This variant regenerates complete plans at fixed intervals, allowing it to adjust to intermediate state observations and potentially improve performance through iterative refinement.
  • Diffusion Forcing [2]: This method represents a variant of diffusion planning that incorporates causal noise scheduling with tokenized trajectories but without explicit search mechanisms. Its default implementation includes replanning capabilities, making it a strong comparative baseline alongside MCTD.
  • MCTD [31]: As our primary baseline, MCTD represents the state-of-the-art in inference-time scaling for diffusion planning. It combines causal noise scheduling with tree-based search, demonstrating exceptional performance across diverse planning tasks. However, our investigation reveals significant computational inefficiencies, particularly in scenarios where its planning capabilities are most valuable.
A.4 Model hyperparameters
While maintaining consistency with the hyperparameter configurations used in prior work [31], but we describe the detailed hyperparameters to improve the reproducibility. We highlight the key parameters specific to our methods while complete implementation details will be available.
Table 7: Diffuser hyperparameters
Hyperparameter Value
Learning Rate $42e-$ 4
EMA Decay $0.995$
Precision in Training/Inference $32$
Batch Size $32$
Max Training Steps $20000$
Planning Horizon Task dependent, discussed in Section A.5
Open Loop Horizon $50$ for Diffuser-Replan, otherwise planning horizon
Guidance Scale $0.1$
Beta Schedule Cosine
Diffusion Model Objective $x_{0}$-prediction
U-Net Depth $4$
Kernel Size 5
The Number of Channels 32, 128, 256
Table 8: Value-Learning Policy Hyperparameters
Hyperparameter Value
Learning Rate $43e-$ 4
Learning Eta $1.0$
Max Q Backup False
Reward Tune cql_antmaze
The Number of Training Epochs $2000$
Gradient Clipping $7.0$
Top-k $1$
Target Steps $10$
Randomness on Data Sampling (p) $0.2$
Table 9: Diffusion Forcing hyperparameters
Hyperparameter Value
Learning Rate $45e-$ 4
Weight Decay $41e-$ 4
Warmup Steps $10000$
Precision in Training $16$-mixed
Precision in Inference $32$
Batch Size $1024$
Max Training Steps $200005$
The Number of Frame Stack $10$
Planning Horizon Task dependent, discussed in Section A.5
Open Loop Horizon $50$
Causal Mask Not Used
Guidance Scale $3$ for medium mazes and $2$ for other mazes
Scheduling Matrix pyramid
Stabilization Level $10$
Beta Schedule Linear
Diffusion Model Objective $x_{0}$-prediction
DDIM Sampling eta $0.0$
Network Size $128$
The Number of Layers $12$
The Number of Attention Heads $4$
The Feedforward Network Dimension $512$
Table 10: MCTDs hyperparameters
Hyperparameter Value
Learning Rate $45e-$ 4
Weight Decay $41e-$ 4
Warmup Steps $10000$
Precision in Training $16$-mixed
Precision in Inference $32$
Batch Size $1024$
Max Training Steps $200005$
The Number of Frame Stack $10$
Planning Horizon Task dependent, discussed in Section A.5
Open Loop Horizon $50$ for Replan otherwise Planning Horizon
Causal Mask Not Used
Scheduling Matrix pyramid
The Maximum Number of Search $500$
Guidance Set Task dependent, discussed in Section A.5
The number of Partial Denoising $20$
The Jumpy Denoising Interval $10$
Stabilization Level $10$
Beta Schedule Linear
Diffusion Model Objective $x_{0}$-prediction
DDIM Sampling eta $0.0$
Network Size $128$
The Number of Layers $12$
The Number of Attention Heads $4$
The Feedforward Network Dimension $512$
Parallelism Degree $200$ for P-MCTD and Fast-MCTD, otherwise $1$
Parallel Search Visitation Count Weight $w$ $1.0$
Subsampling Interval $H$ $5$ for S-MCTD and Fast-MCTD, otherwise $1$
A.5 The evaluation details
Similar to the hyperparameters, many details are consistent with the settings established in prior work [31], while describing key experimental settings below.
A.5.1 Maze navigation with point-mass and ant robots
Guidance set
For pointmaze tasks, $[0,0.1,0.5,1,2]$ is applied and $[0,1,2,3,4,5]$ is set in antmaze evaluation.
Planning horizon
$500$ planning horizon is applied for medium and large sized mazes and $1000$ for giant sized mazes.
Low-level controller
For pointmaze tasks, the heuristic controller [12] is applied and the diffusion-based value-learning policy [28] is utilized for antmaze [4]. For value-learning policy low-level controller acts to the given subgoal which is re-assigned after achieving with $10$ steps interval.
A.5.2 Robot arm cube manipulation
Guidance set
$[1,2,4]$ is applied for each object. MCTDs applies the object-wise guidance, so the size of guidance set is increased as the number of objects increases in the environment.
Planning horizon
$200$ for single cube tasks and $500$ for other tasks.
Low-level control
The value-learning policy [28] is applied as done in antmaze tasks.
Redundant plans between object-wise plans
Following previous work [31], we applied redundant plans to resolve inconsistencies between object-wise plans, such as attempting to move too far from the current position or trying to grasp an object while the robot is already holding another object.
A.5.3 Visual pointmaze
Guidance set.
For visual maze tasks, a guidance scale set of $[0,0.1,0.5,1,2]$ is applied.
Planning horizon.
A fixed planning horizon of 500 steps is used for medium and large-sized mazes.
Low-level controller.
Inverse Dynamics (InvDyna) models were trained following the MCTD framework [31] using offline datasets encoded by a pretrained VAE. These models predict the action $\hat{a}_{t}$ required to transition from the current state $s_{t}$ to the next state $s_{t+1}$, given the partially observable nature of the task. To capture velocity-related information, the InvDyna model conditions its predictions on both the previous state $x_{t-1}$ and the current state $s_{t}$, formalized as:
$\hat{a}_{t}=\mathrm{InvDyna}\bigl{(}(s_{t-1},\,s_{t}),\,s_{t+1}\bigr{)}$ (7)
Guidance in POMDP.
To enhance spatial awareness in partially observable environments, a Position Estimator, pretrained on offline datasets, predicts the agent’s location in the VAE’s latent space. This module provides a coarse but informative positional signal.
Appendix B Additional experimental results
B.1 Leaf parallelization ablations
Table 11: Leaf parallelization ablation study results. Success rates and planning times ($\pm$ std) across PointMaze environment on medium, large giant sized map for Leaf Parallelization ablation.
Success Rate $\uparrow$ (%) Planning Time $\downarrow$ (sec.)
Env. Method medium large giant medium large giant
PointMaze No Leaf Par. (P-MCTD) 100 $\pm$ 0 100 $\pm$ 0 100 $\pm$ 0 8.5 $\pm$ 1.6 7.4 $\pm$ 1.3 12.8 $\pm$ 2.5
Leaf Par. 98 $\pm$ 6 100 $\pm$ 0 100 $\pm$ 0 7.4 $\pm$ 0.9 6.7 $\pm$ 0.6 17.0 $\pm$ 2.8
While P-MCTD implements parallel search through delayed tree updates and search-aware selection, alternative parallelization strategies exist for MCTS [18]. One such approach is leaf parallelization, which expands multiple children from a single selected node in parallel. This technique potentially reduces selection overhead by processing multiple expansions per selection operation.
As shown in Table 11, leaf parallelization yields mixed results across different environment scales. In medium and large mazes, it shows modest efficiency improvements (7.4s vs. 8.5s and 6.7s vs. 7.4s, respectively) with minimal impact on success rates. However, in the giant maze environment, leaf parallelization significantly increases planning time (17.0s vs. 12.8s), demonstrating reduced efficiency precisely where computational optimization is most critical.
This performance degradation in complex environments likely stems from leaf parallelization’s interference with the exploration-exploitation balance maintained by our search-aware selection mechanism, Equation 5. By forcing expansion of multiple children from the same node, leaf parallelization may concentrate computational resources on potentially suboptimal branches, particularly detrimental in larger search spaces where strategic node selection is crucial. These results reinforce our approach of parallelizing across different branches of the search tree rather than within individual branches, especially for complex long-horizon planning scenarios.
B.2 Distillation of Diffusion Planner Ablations
(a) Varying noise levels per subplan
(b) Uniform noise levels per subplan
Figure 6: Distillation challenges in MCTD. Trajectory quality from distilled diffusion models varies significantly based on noise level distribution at training. (a) Using variable noise levels (required for causal scheduling) leads to poor trajectory quality. (b) Using uniform noise levels produces cleaner trajectories but is incompatible with MCTD’s causal scheduling requirements.
Diffusion model distillation represents a promising alternative approach for optimizing denoising processes by training student models to predict the results of multiple denoising steps performed by a teacher model [22]. We conducted a preliminary investigation of this technique as another potential direction for optimizing MCTD.
Following the progressive distillation approach [22], we trained a student model to predict two-step denoised outputs from the teacher model. However, our experiments revealed significant challenges in trajectory quality, as illustrated in Figure 6. Upon analysis, we identified that MCTD’s causal noise scheduling—which applies different noise levels to different subplans—introduces unique complications for distillation approaches. While standard diffusion distillation typically assumes uniform noise levels across the entire generated content, MCTD’s varying per-subplan noise levels create a substantially more complex learning task for the student model.
Our findings suggest that effective distillation for MCTD would require specialized techniques capable of handling these heterogeneous noise distributions. This represents an intriguing avenue for future research, but we leave this as one of the future works.
Appendix C Algorithms
Algorithm 2 Fast Monte Carlo Tree Diffusion
1:procedure Fast-MCTD($root,iterations,parallelism\_degree,jump\_interval$)
2:    for $i=1$ to $iterations$ by $parallelism\_degree$ do
3:        $nodes\_to\_expand\leftarrow[\,]$
4:        $temp\_visit\_counts\leftarrow\{\}$triangleright Track parallel selections
5:        for $j=1$ to $parallelism\_degree$ do $\triangleright$ Search-aware selection
6:           $node\leftarrow root$
7:           while IsFullyExpanded($node$) and not IsLeaf($node$do
8:               $node\leftarrow$ SearchAwareUCT($node,temp\_visit\_counts$)
9:               $temp\_visit\_counts[node]\leftarrow temp\_visit\_counts.get(node,0)+1$
10:           end while
11:           Append($nodes\_to\_expand,node$)
12:        end for
13:        $expansions\leftarrow[\,]$triangleright Parallel expansion phase
14:        for $node\in nodes\_to\_expand$ do
15:           if IsExpandable($node$then
16:               $g_{s}\leftarrow$ SelectMetaAction($node$)
17:               Append($expansions,(node,g_{s})$)
18:           end if
19:        end for
20:        $new\_children\leftarrow$ BatchDenoiseSubplans($expansions,jump\_interval$)
21:        for $(node,child)\in new\_children$ do
22:           AddChild($node,child$)
23:        end for
24:        $simulations\leftarrow[\,]$triangleright Parallel simulation phase
25:        for $(node,child)\in new\_children$ do
26:           $partial\leftarrow$ GetPartialTrajectory($child$)
27:           Append($simulations,(node,child,partial)$)
28:        end for
29:        $rewards\leftarrow$ BatchFastSparseDenoising($simulations,jump\_interval$)
30:        for $(node,child,reward)\in rewards$ do $\triangleright$ Delayed tree update
31:           $current\leftarrow child$
32:           while $current\neq null$ do
33:               $current.visitCount\leftarrow current.visitCount+1$
34:               $current.value\leftarrow\textsc{Max}(current.value,reward)$
35:               UpdateMetaActionSchedule($current,reward$)
36:               $current\leftarrow current.parent$
37:           end while
38:        end for
39:    end for
40:    return BestChild($root$)
41:end procedure